OpenStudy (anonymous):

Coach Johnson is 10 years older than Tim. When Tim started football 3 years ago, Coach Johnson was twice his age. How old is Tim now?

OpenStudy (mathteacher1729):

Quick question -- can you solve equations like: 2x + y = 1 x + y = 2 using algebra?

OpenStudy (anonymous):

C = coach; T = Tim C = T + 10; C-3 = 2(T-3), so C = 2T - 6 + 3 = 2t - 3 2t - 3 = t + 10 t = 13, Tim's age now, so coach is 23; three years ago they were 10 & 20

OpenStudy (mathteacher1729):

(I just made that up off the top of my head, but your answer will help me figure out where to begin helping) :)

OpenStudy (anonymous):

y - 3 = 2x - 3 x + 10 = y y = 2x = x + 10 x = 10, this is Tim's age